OpenStudy (anonymous):

HELPPPPP. The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is 40 in long. What is the length of the other leg? Explain how you found your answer.

OpenStudy (anonymous):

a^2+b^2=c^2, C is always the hypotenuse and a and b are the two other sides. 1600+b=1681

OpenStudy (anonymous):

that b should be b^2

OpenStudy (anonymous):

ok but i dont understand how to explain it

OpenStudy (anonymous):

then you try to get b^2 alone so, 1681-1600=81. so b^2=81

OpenStudy (anonymous):

then you square root both sides and you end up getting b=9

OpenStudy (anonymous):

Do you understand it better now?

OpenStudy (anonymous):

yea. thanks

OpenStudy (anonymous):

okay